### v2

# SGDFuse: SAM-Guided Diffusion for High-Fidelity Infrared and Visible Image Fusion

Abstract
Infrared and visible image fusion (IVIF) aims to combine the thermal radiation information from infrared images with the rich texture details from visible images to enhance perceptual capabilities for downstream visual tasks. However, existing methods often fail to preserve key targets due to a lack of deep semantic understanding of the scene, while the fusion process itself can also introduce artifacts and detail loss, severely compromising both image quality and task performance. To address these issues, this paper proposes SGDFuse, a conditional diffusion model guided by the Segment Anything Model (SAM), to achieve high-fidelity and semantically-aware image fusion. The core of our method is to utilize high-quality semantic masks generated by SAM as explicit priors to guide the optimization of the fusion process via a conditional diffusion model. Specifically, the framework operates in a two-stage process: it first performs a preliminary fusion of multi-modal features, and then utilizes the semantic masks from SAM jointly with the preliminary fused image as a condition to drive the diffusion model’s coarse-to-fine denoising generation. This ensures the fusion process not only has explicit semantic directionality but also guarantees the high fidelity of the final result. Extensive experiments demonstrate that SGDFuse achieves state-of-the-art performance in both subjective and objective evaluations, as well as in its adaptability to downstream tasks, providing a powerful solution to the core challenges in image fusion. The code of SGDFuse is available at https://github.com/boshizhang123/SGDFuse.
keywords:
Image Fusion, Diffusion Model, Segment Anything Model (SAM), Multimodal Information, Image Enhancement
journal: Nuclear Physics B
1 Introduction
Infrared and visible image fusion is an important research direction in the field of computer vision [1]. This technology aims to efficiently combine the thermal radiation information of infrared images with the rich texture details of visible light images, to generate an image that is more comprehensive in information and higher in perceptual clarity [2]. Because infrared images can effectively overcome adverse conditions such as smoke and low light, while visible light images possess the advantages of high resolution and high contrast, the features of the two are highly complementary. This fusion not only significantly improves the environmental perception and target recognition ability in key scenarios such as autonomous driving [3], military reconnaissance, and medical image analysis [4], but also effectively enhances the performance of downstream vision tasks [5]. Therefore, how to effectively achieve fusion to balance the final visual quality and task requirements, is an urgent challenge currently being faced.
In pursuit of this objective, a plethora of techniques have been proposed over the past few decades [6]. Conventional fusion methods are largely predicated on handcrafted feature extraction and fusion rules—such as multi-scale transforms [7], and subspace reconstruction—which offer a degree of interpretability and robustness. However, these approaches face inherent limitations in modeling the high-order semantic relationships and non-linear complementarities between modalities, thereby restricting their efficacy in complex scenes [8]. With the advent and rapid proliferation of Deep Learning (DL), the research landscape for IVIF has undergone a paradigm shift. This has yielded substantial advancements in feature extraction, fusion, and reconstruction, rapidly establishing the field as a prominent area of research. For instance, Li et al. introduced DenseFuse [9], which leverages an encoder-decoder architecture to extract multi-level features; PIAFusion [10], proposed by Tang et al., employs a progressive fusion strategy to handle extreme variations in illumination; Text-DiFuse [11] achieves controllable fusion and enhanced robustness against degradation by integrating text modulation with diffusion-based generation.
Figure 1: Comparison of different fusion methods on object detection and semantic segmentation tasks. Traditional methods perform well in terms of visual quality but are limited in downstream tasks, while the proposed method achieves better task performance while maintaining image quality.
Despite advancements in feature extraction and fusion strategies, inherent design limitations in existing methods lead to a core problem: a ’semantic blindness’. For instance, methods based on CNNs are constrained by their local receptive fields, an architectural limitation that hinders their ability to comprehend global scene structures, often resulting in blurred boundaries for large-scale targets or inconsistent background luminance. On the other hand, while GAN-based approaches can produce visually sharp images, their training relies on pixel-level fidelity constraints (e.g., intensity and gradient losses), which struggle to capture complex cross-modal semantic relationships. This not only causes training instability but also frequently leads to the erroneous suppression of salient thermal targets in infrared images. These specific design flaws point to a fundamental issue: the models are unable to semantically differentiate between foreground targets and background textures, thus failing to perform selective information enhancement. This ”can’t see the forest for the trees” paradigm directly causes a cascade of problems, including blurred target boundaries, loss of critical structures, and erroneous suppression of thermal features. Ultimately, this semantic deficiency is the fundamental bottleneck constraining the efficacy of fused images in downstream tasks like object detection and semantic segmentation. As illustrated in Fig. 1, the failure of existing models to preserve task-critical target regions severely undermines the practical utility of fusion technology.
Recent advancements in large-scale vision models offer new promise [12]. The Segment Anything Model (SAM) [13] represents a key breakthrough owing to its outstanding semantic segmentation capabilities. The high-quality semantic masks it generates can provide multi-modal images with rich semantic priors (as shown in Fig. 2). This wealth of semantic information can significantly alleviate the ”semantic blindness” inherent in existing methods. It guides the model to more effectively preserve the thermal radiation features and visible details of critical targets, thereby achieving a more semantically consistent fused representation. On the other hand, image fusion, as an image generation task, is typically accompanied by processes of information compression and modal reconstruction [14]. Although this process can enhance visual consistency, it often leads to a degradation in detail quality, manifesting as blurred textures, indistinct boundaries, and visual artifacts. Such degradation not only diminishes the perceptual quality of the fused image but also significantly impairs the performance of downstream tasks that rely on structural information, such as object detection and semantic segmentation. Diffusion models, by virtue of their
Figure 2: Visible and infrared images along with their corresponding SAM semantic masks. These semantic cues are used to guide the diffusion model optimization in the second stage.
iterative denoising generation mechanism, can stably produce high-fidelity, artifact-free images, making them exceptionally well-suited for the demands of IVIF, namely detail restoration and modal balance. More importantly, when integrated with high-quality semantic priors, diffusion models can perform semantically-aware conditional generation. This enables the fusion process to be more targeted in its representation of structures, textures, and target regions, significantly enhancing its adaptability for downstream tasks.
Building upon the aforementioned insights, we propose the SAM-Guided Diffusion Fusion Network (SGDFuse), which employs a two-stage training framework to achieve synergistic optimization of structural semantic understanding and high-fidelity image generation. The first stage focuses on multi-modal feature extraction and fusion. We design a Multi-Scale Feature Enhancement Module (MSFEM) and a Transformer Block to model the local structures of the infrared image and the global semantics of the visible image, respectively. A preliminary fused image is then generated through the effective alignment of these cross-modal features. The second stage is oriented towards task-oriented optimization and high-fidelity reconstruction. Here, we construct a fused image optimization mechanism based on a conditional diffusion model, concentrating on structural detail restoration and task-relevant semantic enhancement. Specifically, we leverage SAM to generate rich semantic prior masks for both the infrared and visible images. These masks are concatenated with the preliminary fused image to form a five-channel, task-aware input. This input guides the conditional diffusion process to progressively reconstruct the image structure and enhance semantic details. Concurrently, we design a Hierarchical Feature Aggregation Head (HFAH) to aggregate deep semantic information from multiple scales, further improving the fused image’s representation of target structures and regional boundaries. Through the organic integration of semantic guidance and high-fidelity reconstruction, our method establishes an effective synergy between structural understanding and detail generation. This significantly boosts the semantic integrity, visual quality, and adaptability for downstream tasks of the resulting fused image.
Our main contributions are summarized as follows:
  • 1.
    Semantic-Aware Fusion: By utilizing high-quality semantic masks from SAM as explicit priors, our method effectively overcomes the ”semantic blindness” of conventional approaches. This results in a superior preservation and enhancement of key semantic information, like infrared thermal targets and visible textures, boosting the semantic integrity of the fused image.
  • 2.
    High-Fidelity Image Optimization Mechanism: We pioneer the use of a conditional diffusion model to perform high-fidelity optimization on the fused image. By virtue of their robust training stability and excellent generative power, diffusion models enable a high-precision reconstruction of the preliminary fused information, effectively minimizing artifacts and ensuring the final representation maintains maximum fidelity under semantic guidance.
  • 3.
    Two-Stage Task-Oriented Fusion Framework: We propose a framework that innovatively unifies structural understanding and high-quality generation by synergistically integrating multi-modal feature fusion with task-aware, diffusion-based optimization. This approach, guided by semantic priors, effectively boosts the fused image’s performance on downstream tasks and its representation of regional structures.
2 Related work
2.1 Deep Multimodal Image Fusion
Recent advances in deep learning have significantly propelled multimodal image fusion (IVIF), which integrates information from different sensors or viewpoints to produce clearer, higher-quality images [15]. Leveraging deep neural networks’ strong feature learning and nonlinear modeling abilities, various data-driven infrared and visible image fusion methods have been developed, including those based on autoencoders, CNNs, GANs, and Transformers.
Autoencoder (AE)-based methods, such as DenseFuse [9], utilize an encoder-decoder architecture for multi-scale feature extraction and fusion. While advanced variants like SEDRFuse [16] introduce refinements using residual and attention mechanisms, their fundamental reliance on pixel-level loss restricts the capture of deep semantics, often leading to a loss of critical contextual information.
Methods based on Convolutional Neural Networks (CNNs) enhance feature extraction. Jiang et al. [17] ingeniously designed a harmonized representation space for multi-modality representation and achieve one of the best infrared and visible registration results. Li et al. [18] proposed a novel multimodal medical image fusion framework that leverages semantic-level information to offer clinicians a more efficient way of observing lesions across different imaging modalities. However, their fundamental reliance on local receptive fields limits the ability to model global context and long-range dependencies, thereby compromising the semantic consistency of the fused image.
Transformer-based methods provide a new fusion paradigm, leveraging long-range dependency modeling for superior cross-domain integration. Various works have employed Transformers to enhance inter-modal dependencies, suppress redundancy, and integrate complementary features [19]. Building on this foundation, research has extended into more specialized and cutting-edge applications. Li et al. [20] proposed a U-Net-based framework, pioneering the exploration of adversarial robustness in fusion tasks. To achieve precise semantic guidance, Liu et al. [21] pioneered the application of prompt learning to infrared and visible image fusion, with achieving sota performance.
Generative Adversarial Network (GAN)-based approaches introduced an unsupervised adversarial learning paradigm to image fusion. FusionGAN [22] uses adversarial training between generator and discriminator. AttentionFGAN [23] adds multi-scale attention to separately emphasize infrared foreground and visible background. However, despite their excellent feature preservation, these methods often suffer from training instability. Furthermore, the generator’s reliance on low-level fidelity constraints (e.g., gradient, intensity) hinders the modeling of complex inter-modal distributions in the latent space, thereby compromising high-level semantic integrity.
By contrast, our method introduces semantic prior masks and a high-fidelity generation mechanism to produce fused images that are both semantically consistent and rich in detail, significantly enhancing support for downstream tasks.
2.2 SAM and Its Applications
The Segment Anything Model (SAM) [13], as a powerful pretrained model, has become a current research hotspot in the field of image processing due to its zero-shot generalization capability in image segmentation and object detection tasks [24]. SAM can provide high-precision semantic masks across various vision tasks and has been widely applied in low-level vision tasks such as dehazing [25], and low-light image enhancement [26], as well as high-level tasks like object detection [27], demonstrating outstanding performance. By deeply understanding image content, SAM effectively captures fine details and accurately delineates target regions. Therefore, SAM is particularly suitable for infrared and visible image fusion tasks, we leverage SAM-generated segmentation masks as explicit, high-level semantic priors to guide the fusion process.
2.3 Diffusion Models
Diffusion models have emerged as a powerful class of deep generative models, achieving state-of-the-art results in image generation, restoration, super-resolution, and image-to-image translation tasks [5], [28]. Beyond generative applications, their learned representations have also shown potential in discriminative tasks such as classification, segmentation, and object detection. A diffusion model consists of two stages: a forward process that gradually adds Gaussian noise to the input data, and a reverse process that iteratively denoises the corrupted data by predicting and removing noise, ultimately recovering high-quality samples [29].
The advantages of diffusion models lie in their ability to generate high-quality and diverse samples while effectively capturing deep underlying features of the data, which leads to outstanding performance in image generation tasks. With the continuous advancement of diffusion models across various fields, particularly in image generation, they have begun to surpass the long-standing dominance of generative adversarial networks (GANs). In the context of infrared and visible image fusion, diffusion models provide powerful support for generating high-quality fused images.
Therefore, This paper explores a diffusion model-based approach for infrared and visible image fusion, aiming to achieve superior fusion results through this generative paradigm, thereby enhancing both task performance and image quality.
3 Methodology
3.1 Motivation
Existing IVIF methods face a core bottleneck in semantic understanding, as they treat fusion as a mathematical reorganization of low-level visual features (e.g., pixel intensity, gradients). This failure to semantically distinguish between critical targets and background textures acts as a fundamental obstacle, severely limiting the potential of the fused images in downstream tasks like object detection and semantic segmentation.
Our primary and core motivation is to inject explicit semantic understanding capabilities into the fusion process. We are no longer satisfied with the indiscriminate processing of low-level features by existing methods, but rather seek a novel mechanism that enables the model to truly ”understand” the image and distinguish ”target” from ”background”. For this, we introduce SAM. The high-quality semantic masks it generates can provide the fusion process with unprecedented, strong semantic priors concerning scene structure and object boundaries, which becomes our ideal tool for solving the core problem of ”semantic blindness”. However, high-level semantic guidance alone is insufficient. To accurately translate this precise semantic intent—for instance, ’to preserve the thermal features of a target while enhancing the texture of the background’—into a high-quality, pixel-level output, one must rely on an equally powerful high-fidelity generative framework. Traditional reconstruction methods are prone to generating artifacts and detail loss when handling such complex conditions, which severely undermines the efficacy of semantic guidance. Therefore, our second motivation is to employ a Diffusion Model, renowned for its superior generative capabilities, as the backbone for image reconstruction. This ensures that the high-level semantic commands
Figure 3: Overall architecture of the SGDFuse framework. The framework consists of two stages: the first stage performs multi-modal feature extraction and generates an initial fused image; the second stage incorporates semantic prompts and a conditional diffusion network to refine structural details and enhance semantic consistency of the fused image.
can be rendered losslessly and with maximum fidelity.
The core of our paradigm is the Deep Coupling of semantic understanding and image generation, rather than treating them as isolated steps. We leverage high-quality semantic masks from SAM as strong conditional priors to guide the iterative denoising of a conditional diffusion model. This endows the generation process with Precise Semantic Directivity, enabling the model to intelligently preserve thermal targets while reconstructing texture details based on semantic importance. Ultimately, SGDFuse unifies the contradictory objectives of semantic integrity and visual quality, producing fused images that are both visually compelling and highly effective for downstream tasks.
3.2 Network Architecture
In this section, we provide a detailed description of the proposed image fusion framework, SGDFuse. This fusion framework adopts a two-stage architecture, as illustrated in Fig. 3, aiming to generate high-quality fused images with structural fidelity and semantic consistency.
In Stage I, SGDFuse generates an initial fused image via a tailored multimodal feature extraction and fusion scheme. The infrared image (IR) is processed by the Multi-Scale Feature Enhancement Module (MSFEM), which uses parallel convolutional branches and channel attention to capture thermal boundaries and enhance structural cues at multiple receptive fields. Meanwhile, the visible image (VIS) is encoded by a Transformer Block (TB) with multi-head self-attention to extract global context and fine-grained texture. The extracted features are then dynamically aligned and selectively fused through a cross-attention-based interaction pathway, yielding an initial fused image that combines salient IR targets with VIS texture details.
In Stage II, SGDFuse applies a task-aware conditional diffusion mechanism to refine structure and enhance semantic consistency. Specifically, SAM generates high-quality semantic masks from the IR and VIS images, which are concatenated with the initial fused image to form a five-channel input guiding the diffusion process. A denoising decoder progressively reconstructs a high-fidelity fused image via reverse diffusion. Additionally, a Hierarchical Feature Aggregation Head (HFAH) integrates semantic features across diffusion stages, strengthening edge, structure, and region consistency, thereby improving generalization in downstream tasks.
3.3 Multi-Scale Feature Enhancement Module (MSFEM)
To better capture target edges and thermal regions in infrared images, SGDFuse uses a Multi-Scale Feature Enhancement Module (MSFEM) in stage one to enhance structural details across scales. Since infrared images have salient targets but blurred edges and missing details, MSFEM applies multi-scale modeling with attention to strengthen key region representations.
As shown in Fig. 3(c), the MSFEM employs a multi-branch parallel structure, where each branch utilizes convolutional operations with different receptive fields (e.g., $\times1$, $\times3$, $\times5$, $\times7$) to capture structural features ranging from local to medium scales. Specifically, given an input image, feature representations are first extracted separately by four sets of convolutional kernels:
$F_{i}=Conv_{j\times j}(F_{ir})(i=1,2,3,4;j=1,3,5,7).$ (1)
Then, the features from the three large-scale branches $F_{2}$, $F_{3}$ and $F_{4}$ are concatenated along the channel dimension to obtain a fused multi-scale representation:
$F_{ms}=Concat(F_{2},F_{3},F_{4}).$ (2)
To enhance feature representation, the concatenated feature $F_{ms}$ is sequentially processed through modules $DWConv_{3\times 3}$, $Conv_{1\times 1}$ and $DWConv_{3\times 3}$ for feature enhancement and compression. This process can be expressed as follows:
$F_{enh}=DWConv_{3\times 3}(Conv_{1\times 1}(DWConv_{3\times 3}(F_{ms}))).$ (3)
Meanwhile, the $\times1$ convolution branch $F_{1}$ preserves shallow-level details and linear response capabilities, providing complementary information to the main features. Finally, features $F_{ms}$ and $F_{1}$ are concatenated and fused:
$F_{cat}=Concat(F_{enh},F_{1}).$ (4)
To compress the channel dimension and unify the scale, the fused features $F_{cat}$ are passed through a $Conv_{1\times 1}$ for channel reduction, followed by a Sigmoid activation function to obtain normalized response weights. Finally, MSFEM employs a residual connection by element-wise addition of the normalized weights with the original input features , producing the module’s final output. This process can be expressed as follows:
$F_{out}=F_{ir}+\sigma(Conv_{1\times 1}(F_{cat})).$ (5)
The residual mechanism aids feature propagation, improves training stability, and enhances nonlinear representation. Ablation studies show that stacking MSFEM three times yields optimal fusion, confirming its effectiveness in multi-scale structural modeling and infrared feature enhancement.
3.4 Semantic-Aware Fusion Optimization Driven by Conditional Diffusion
To enhance structural refinement and semantic consistency of the fused image, this work introduces an optimization mechanism based on a conditional diffusion model in the second stage. Unlike existing methods that directly concatenate the original infrared and visible images, our approach concatenates the initial fused image $F_{1}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$ generated in the first stage with two semantic masks $M_{ir}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 1}$ and $M_{vis}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 1}$ extracted from the infrared and visible images using the Segment Anything Model (SAM), forming a five-channel task-aware input. We employ the Denoising Diffusion Probabilistic Model (DDPM) [30] to model the diffusion generation process of this five-channel input in the latent space. The forward process gradually adds Gaussian noise over T time steps to produce a perturbed data distribution, while the reverse process progressively denoises the input under the guidance of semantic masks to reconstruct a fused image with clear structure and semantic consistency. By modeling the joint distribution of the fused image and semantic priors, the diffusion model learns a semantic-aware generative pathway, thereby further improving fusion quality and adaptability to downstream tasks.
Figure 4: Architecture of the denoising network and the Hierarchical Feature Aggregation Head (HFAH).
3.4.1 Forward Diffusion Process
The diffusion process consists of a forward perturbation process (Forward Process) and a reverse generation process (Reverse Process), whose core idea is derived from non-equilibrium thermodynamics theory [31] and is systematically modeled in DDPM [30]. In the forward diffusion process, the five-channel input $\mathrm{I}_{0}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 5}$, formed by concatenating the initial fused image and semantic masks, is gradually perturbed by Gaussian noise through a Markov chain of T time steps, causing its distribution to progressively approach a standard Gaussian distribution. Specifically, the noisy image $I_{t}$ at the t-th time step can be represented by the following conditional distribution:
$P(I_{t}|I_{t-1})=\mathcal{N}(I_{t};\sqrt{\alpha_{t}}I_{t-1},(1-\alpha_{t})\mathbf{I}),$ (6)
where $\alpha_{\mathrm{t}}\in(0,1)$ denotes the variance scheduling factor at the t-th time step, $\mathbf{I}$ represents the identity covariance matrix, and the Gaussian noise follows a standard normal distribution.
Furthermore, to improve sampling efficiency, the model supports directly generating the corresponding noisy image $I_{t}$ at any arbitrary time step t based on the original input $I_{0}$. The calculation is formulated as follows:
$I_{t}=\sqrt{\bar{\alpha}_{t}}I_{0}+\sqrt{1-\bar{\alpha}_{t}}\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,\mathbf{I}),$ (7)
where $\overline{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ denotes the cumulative noise retention factor up to time step t, and $\varepsilon$ is a noise variable sampled from the standard Gaussian distribution.
In this study, the five-channel input is constructed by concatenating the three-channel preliminary fused image $F_{1}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$ with two semantic masks $M_{ir}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 1}$ and $M_{vis}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 1}$ generated by the Segment Anything Model (SAM):
$I_{0}=\mathrm{Concat}(F_{1},M_{ir},M_{vis})\in\mathbb{R}^{H\times W\times 5}.$ (8)
This design enables the diffusion process to model not only the visual appearance of the image but also incorporate task-relevant semantic guidance, thereby effectively enhancing structural preservation and semantic consistency during the subsequent reverse generation process.
3.4.2 Reverse Diffusion Process
During the reverse diffusion process, the diffusion model employs a parameterized neural network to progressively denoise the noisy image at each time step, ultimately restoring a high-quality fused image relevant to the task. Specifically, at each time step $t$, the network estimates the current noisy image $I_{t}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 5}$ to obtain the image at the previous time step $I_{t-1}$. This process can be modeled as a conditional Gaussian distribution:
$Q(I_{t-1}|I_{t})=\mathcal{N}(I_{t-1};\mu_{\theta}(I_{t},t),\sigma_{t}^{2}\mathbf{I}),$ (9)
where $\mu_{\theta}(I_{t},t)$ denotes the mean predicted by the neural network at the current time step $t$, and $\sigma_{t}^{2}$ represents the noise variance at this time step, which is expressed as follows:
$\sigma_{t}^{2}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t},$ (10)
where $\beta_{t}=1-\alpha_{t}$ denotes the noise level added at each step, and $\bar{\alpha}_{t}$ is the cumulative noise retention factor up to step t. The mean term $\mu_{\theta}(I_{t},t)$ of the conditional distribution can be expressed as follows:
$\mu_{\theta}(I_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(I_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\cdot\varepsilon_{\theta}(I_{t},t)\right),$ (11)
where $\varepsilon_{\theta}(I_{t},t)$ represents the noise predicted by the denoising network (U-Net) for the current image $I_{t}$ at time step t.
In the proposed method, the input image $I_{t}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 5}$ consists of the three-channel initial fused image concatenated with two semantic masks. Therefore, the reverse diffusion process not only restores detailed structures but also enhances semantic consistency under the guidance of semantic prompts. By iteratively performing this process, the diffusion model progressively reconstructs a high-fidelity and semantically coherent final fused image from standard Gaussian noise.
3.4.3 Loss Function of Diffusion Process
In this work, we adopt the simplified training objective consistent with DDPM [30], which aims to minimize the mean squared error between the noise predicted by the model and the actual added Gaussian noise. The loss function is expressed as follows:
$\mathcal{L}_{\mathrm{diff}}=\left\|\gamma-\varepsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}I_{0}+\sqrt{1-\bar{\alpha}_{t}}\gamma,t\right)\right\|_{2}^{2},$ (12)
where $I_{t}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 5}$ denotes the original five-channel input image formed by concatenating the preliminary fused image and semantic masks; $\gamma\in\mathrm{N}(0,\mathrm{I})$ represents the noise sampled from a standard normal distribution; $\bar{\alpha}_{t}$} is the cumulative retention coefficient up to time step $t$ and $\varepsilon_{\theta}$ is the parameterized denoising network. By minimizing this loss, the model learns to effectively recover the clean image from noisy inputs at any time step, enabling a semantic-aware denoising generation process.
3.4.4 Structure of the Denoising Network
In the second stage of our framework, to accurately estimate the Gaussian noise added during the forward diffusion process, we design a denoising network $\varepsilon_{\theta}(\cdot,\cdot)$ based on a U-Net architecture. Inspired by the design principles of SR3 [31], the proposed network exhibits strong image restoration capability and multi-scale representation learning, enabling effective modeling of the latent structure within the semantically guided five-channel input. Specifically, the denoising network comprises three main components: a contracting path (consisting of five convolutional layers) that progressively downsamples the input to extract deep semantic and structural features; an expanding path (with five corresponding convolutional layers) that incrementally restores spatial resolution; and a diffusion head that outputs noise predictions matching the input dimensions. Leveraging the five-channel input formed by concatenating the preliminary fused image and semantic masks, the network is capable of capturing both low-level texture and edge structures as well as high-level semantic cues, thereby facilitating semantic-aware high-fidelity image reconstruction.
3.5 Hierarchical Feature Aggregation Head (HFAH)
To further enhance feature representation in the decoding stage, we introduce a Hierarchical Feature Aggregation Head (HFAH) into the decoder path of the denoising network. HFAH integrates multi-level decoded features and incorporates a spatial attention mechanism to jointly optimize structural detail and semantic consistency. Specifically, all weighted hierarchical features are concatenated and passed through a fusion head to generate the final fused image $I_{f}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$, which exhibits clearer structural boundaries and improved detail preservation. The fusion head consists of multiple $\times3$ convolutional layers designed to map the high-dimensional, multi-scale diffusion features into a final three-channel fused image. A Tanh activation function is applied at the output to enhance texture continuity and fine detail expression. Benefiting from HFAH’s precise modeling of structural edges and semantic regions, this module significantly improves the boundary fidelity and fine-grained reconstruction quality of the final fused image.
3.6 Loss Function of Training Process
To achieve semantically consistent and structurally clear multimodal image fusion, this paper designs task-specific loss functions under the two-stage training framework of SGDFuse. Specifically, dedicated loss objectives are formulated for the structural representation of the preliminary fused image in the first stage, and for the high-quality generation of the final fused image in the second stage.
3.6.1 Stage I: Supervision of Preliminary Fusion
The first stage aims to achieve modal alignment and information fusion between infrared and visible images, producing a preliminary fused image $F_{1}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$. Gradient Loss and Intensity Loss are employed in this stage to ensure structural consistency with the visible image and preserve thermal information from the infrared image, defined as:
$L_{grad}^{1}=\frac{1}{HW}\parallel\nabla F_{1}-\nabla I_{vis}\parallel,$ (13)
$L_{int}^{1}=\frac{1}{HW}\parallel F_{1}-I_{ir}\parallel,$ (14)
where $\nabla$ denotes the image gradient operator used to extract edge structures, H and W are the height and width of the image.
The overall loss for the first stage is defined as:
$L_{stage1}=L_{int}^{1}+L_{grad}^{2},$ (15)
3.6.2 Stage II: Diffusion Optimization Training
In the second stage, the preliminary fused image is concatenated with the infrared and visible semantic masks obtained from SAM to form a five-channel input, which guides the conditional diffusion model to generate the final fused image $I_{f}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$. To further enhance structural fidelity and semantic consistency, the fusion process is supervised using the following loss functions.
Mask-guided Intensity Loss: To enhance the luminance consistency and thermal response representation of the fused image in semantically salient regions, a region-weighted intensity loss is introduced. Let the infrared and visible semantic masks generated by SAM be denoted as $M_{ir}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 1}$ and $M_{vis}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 1}$, respectively. A joint region mask $M\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 1}$ is constructed as follows:
$\mathrm{M=max(M_{ir},M_{vis})},$ (16)
$L_{int}^{mask}=\frac{1}{HW}\parallel M\cdot(I_{f}-max(I_{ir},I_{vis}))\parallel_{1},$ (17)
where $I_{f}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$ denotes the final fused image.
Mask-guided Gradient Loss: To further enhance edge clarity within salient regions, a region-weighted gradient loss is introduced. This loss guides the fused image to align with the clearer edges from either the infrared or visible image within the salient areas. It is defined as follows:
$L_{grad}^{mask}=\frac{1}{HW}\parallel M\cdot(\nabla I_{f}-max(\nabla I_{ir},\nabla I_{vis}))\parallel_{1},$ (18)
where $\nabla$ denotes the image gradient operator used to extract edge structures; the loss is applied within the masked regions to guide the network to focus on optimizing the structural edges of the target areas.
Finally, the total loss for the second stage is defined as:
$L_{stage2}=\lambda_{1}\cdot L_{int}^{mask}+\lambda_{2}\cdot L_{grad}^{mask},$ (19)
where $\lambda_{1}$ and $\lambda_{2}$ are weighting coefficients that control the relative importance of regional intensity alignment and edge alignment. Based on experimental validation, the optimal values are determined as $\lambda_{1}$=1.5 and $\lambda_{2}$=1.
Table 1: Quantitative comparison on MSRS and M3FD datasets, with red representing the best results, blue representing the second best results.
Method MSRS M3FD
EN SD SF MI SCD VIF $Q^{abf}$ EN SD SF MI SCD VIF $Q^{abf}$
U2Fusion[1] 5.37 25.52 9.07 1.4 1.24 0.51 0.42 6.28 26.26 8.36 1.79 1.22 0.68 0.48
TarDAL[32] 5.28 25.22 5.98 1.49 0.71 0.44 0.18 6.81 38.57 6.78 2.27 1.21 0.60 0.31
PIAFusion [10] 6.64 45.34 12.12 2.75 1.7 1.02 0.67 6.87 34.01 12.49 2.61 1.43 0.94 0.60
IGNet[27] 6.12 33.91 9.66 1.29 1.56 0.69 0.47 6.94 39.93 12.17 1.45 1.54 0.74 0.45
CDDFuse [33] 6.70 43.38 11.55 3.43 1.62 1.05 0.69 6.85 37.47 12.87 2.27 1.60 0.90 0.61
Dif-fusion[5] 6.47 37.69 10.86 1.78 1.47 0.82 0.53 6.59 32.71 11.41 2.01 1.19 0.48 0.42
PFCFuse [34] 6.69 43.12 11.60 2.65 1.63 0.92 0.61 6.85 37.62 13.29 2.15 1.55 0.69 0.51
SHIP [24] 6.43 41.13 11.81 2.86 1.51 1.04 0.66 6.63 32.46 13.45 2.91 1.11 0.86 0.62
EMMA [28] 6.71 44.13 11.56 2.94 1.63 0.97 0.64 6.93 38.31 12.84 2.79 1.49 0.66 0.58
Text-Difuse[11] 6.65 44.96 12.97 2.89 1.45 0.84 0.69 6.91 37.78 13.08 2.91 1.63 0.95 0.61
DAFusion [15] 6.69 45.02 12.84 2.56 1.62 0.91 0.68 6.94 37.78 13.08 2.91 1.63 0.95 0.63
MaeFuse [6] 6.74 45.19 12.96 2.97 1.69 1.06 0.72 6.96 37.98 13.24 2.97 1.68 1.00 0.66
SAGE [12] 6.71 45.04 13.17 2.91 1.71 1.03 0.70 6.87 37.86 13.15 2.94 1.65 0.97 0.64
SGDFuse(Ours) 6.81 45.28 13.27 2.99 1.73 1.08 0.74 6.99 38.02 13.27 2.99 1.69 1.01 0.69
Table 2: Quantitative comparison on LLVIP and RoadScene datasets, with red representing the best results, blue representing the second best results.
Method LLVIP RoadScene
EN SD SF MI SCD VIF $Q^{abf}$ EN SD SF MI SCD VIF $Q^{abf}$
U2Fusion[1] 5.33 18.05 6.73 1.52 0.61 0.45 0.23 6.59 38.11 10.73 1.75 1.68 0.53 0.46
TarDAL[32] 6.01 27.44 7.02 1.63 0.83 0.55 0.22 7.15 47.68 10.10 2.18 1.56 0.60 0.40
PIAFusion [10] 7.03 42.88 16.10 2.26 1.54 0.91 0.65 6.99 44.23 11.51 2.47 1.58 0.69 0.46
IGNet[27] 6.88 42.95 13.95 1.44 1.53 0.78 0.50 7.23 42.47 9.72 1.38 1.50 0.65 0.47
CDDFuse [33] 7.01 42.47 14.52 2.89 1.56 0.93 0.63 7.44 54.58 16.29 2.30 1.80 0.69 0.52
Dif-fusion[5] 7.06 43.30 14.48 2.02 1.47 0.72 0.52 7.10 44.90 10.11 2.19 1.52 0.52 0.41
PFCFuse [34] 6.87 40.62 15.26 2.62 1.52 0.90 0.63 7.41 50.37 13.81 2.37 1.78 0.74 0.55
SHIP [24] 6.89 41.71 15.84 2.63 1.39 0.92 0.66 7.15 44.67 14.59 2.83 1.41 0.71 0.58
EMMA [28] 6.74 42.68 14.75 2.41 1.37 0.87 0.59 7.14 48.73 10.78 2.76 1.49 0.68 0.54
Text-Difuse[11] 6.99 42.94 15.73 2.58 1.54 0.90 0.65 7.21 50.29 9.49 1.74 1.24 0.77 0.37
DAFusion [15] 7.04 43.13 16.01 2.64 1.57 0.89 0.66 7.19 50.37 11.59 1.75 1.69 0.78 0.57
MaeFuse [6] 7.08 43.24 16.13 2.72 1.62 0.96 0.69 7.26 51.64 14.96 2.72 1.72 0.79 0.60
SAGE [12] 7.09 43.19 16.09 2.67 1.59 0.93 0.67 7.21 51.24 14.73 2.59 1.79 0.75 0.59
SGDFuse(Ours) 7.12 43.35 16.15 2.79 1.65 0.99 0.71 7.29 51.79 15.68 2.89 1.84 0.82 0.62
Figure 5: Qualitative comparisons between our method and thirteen state-of-the-art fusion approaches on the MSRS dataset.
Figure 6: Qualitative comparisons between our method and thirteen state-of-the-art fusion approaches on the M3FD dataset.
4 EXPERIMENTS
4.1 Experimental Settings
4.1.1 Datasets
[…]
4.1.2 Evaluation Metrics
We employ seven commonly used metrics for quantitative evaluation. Specifically, these include Entropy (EN), Standard Deviation (SD), Spatial Frequency (SF), Mutual Information (MI), Spectral Color Distortion (SCD), and Visual Information Fidelity (VIF). Higher values of these metrics indicate better image quality.
4.1.3 Training Details
The proposed model is trained on the MSRS dataset, comprising 1,083 visible-infrared image pairs for training and 361 pairs for testing. Following [5], we crop $\times160$ image patches during training to enhance data diversity. Inspired by [36], multi-channel diffusion features are extracted across multiple time steps to enrich representations. The fusion network is trained for 200 epochs using the Adam optimizer [37] with a learning rate of 1e-4 and batch size of 24. During testing, the model processes input images at their original resolution, outputting a three-channel color fused image suitable for subsequent qualitative and quantitative evaluations. The entire model is implemented using the PyTorch framework, and all experiments are conducted on a workstation equipped with an NVIDIA RTX 4090 GPU.
4.2 Generalization Comparison Experiments
To evaluate the generalization capability of our network, we compare it against the following representative image fusion methods: U2Fusion [1], TarDAL [38], PIAFusion [10], IGNet [27], CDDFuse [32], Dif-fusion [5], PFCFuse [34], SHIP [24], EMMA [28], Text-DiFuse [11], DAFusion [15], MaeFuse [6], and SAGE [12].
4.2.1 Results and Analysis on the MSRS Dataset
As shown in Table I, our method achieves the best performance across five key metrics: EN, SD, SF, VIF, and $Q^{abf}$. The higher EN and MI values indicate better information preservation from both modalities, while superior SD and SF reflect enhanced contrast and edge sharpness. Notably, the highest $Q^{abf}$ further confirms the effectiveness of our method in maintaining structural consistency and perceptual quality.
Fig. 5 shows qualitative comparisons under diverse conditions, including varying illumination, multiple targets, and day/night scenes. While U2Fusion and PLAFusion exhibit blurred details and brightness inconsistency, IGNet, CDDFuse, and MaeFuse partially enhance thermal targets but often lack texture fidelity or introduce artifacts. In contrast, our method effectively preserves thermal saliency, enhances structural clarity, and maintains global luminance balance, achieving superior perceptual quality and semantic consistency.
4.2.2 Results and Analysis on the M3FD Dataset
As shown in Table I, our method achieves the best performance across five key metrics: EN, MI, SCD, VIF, and $Q^{abf}$. The high EN (6.99) and MI scores indicate richer information content and stronger source correlation. The superior SCD reflects better structural preservation, while the leading VIF and $Q^{abf}$ scores demonstrate enhanced visual quality and edge consistency.
Fig. 6 shows qualitative comparisons on challenging scenes from the M3FD dataset, including nighttime roads, uneven lighting, low-contrast IR targets, and complex textures. Methods like U2Fusion and TarDAL suffer from blurred structures and missing textures, while IGNet, PLAFusion, and CDDFuse offer partial improvements but still lack edge continuity and luminance balance. MaeFuse and SAGE improve global coherence but exhibit limitations in detail preservation. In contrast, our method consistently restores thermal contours and visible textures with natural brightness and clear contrast. Especially in scenes with weak IR signals and cluttered backgrounds, SGDFuse achieves accurate foreground–background separation, highlighting its superiority in structural refinement and semantic consistency.
4.2.3 Results and Analysis on the LLVIP Dataset
As shown in Table II, our proposed method attains the highest scores across all evaluation metrics
Figure 7: Qualitative comparisons between our method and thirteen state-of-the-art fusion approaches on the LLVIP dataset.
Figure 8: Qualitative comparisons between our method and thirteen state-of-the-art fusion approaches on the RoadScene dataset.
except MI, where it ranks second, demonstrating strong competitiveness. These results highlight SGDFuse’s clear superiority in preserving information content, texture details, edge sharpness, and perceptual quality. Notably, leading performance in EN, SD, and SF indicates richer information, enhanced contrast, and improved clarity in the fused images. Furthermore, the superior $Q^{abf}$ and VIF scores confirm that SGDFuse achieves better perceptual consistency and edge alignment, closely aligning with human visual preferences.
The LLVIP dataset mainly contains nighttime scenes with challenges such as strong occlusion, weak textures, and extremely low illumination, which place higher demands on fusion robustness. Fig. 7 presents visual comparisons under these conditions. Methods like U2Fusion and TarDAL fail to enhance thermal saliency and suffer from blurred details and suppressed backgrounds. SHIP and SAGE improve thermal contrast but struggle with texture reconstruction and luminance balance. In contrast, our method preserves IR thermal responses while effectively integrating VIS textures and structures, producing results with sharper edges, consistent brightness, and stronger global contrast. Even under severe low-light conditions, SGDFuse demonstrates high stability and perceptual quality, confirming its robustness and practical applicability.
Figure 9: Comparison of different fusion methods in terms of parameter count and computational complexity.
4.2.4 Results and Analysis on the RoadScene Dataset
As shown in Table II, our method attains the best performance on four key metrics: MI, SCD, VIF, and $Q^{abf}$. The leading MI score evidences effective preservation of complementary information from infrared and visible inputs. Superior SCD highlights strong structural and contour fidelity. Additionally, top VIF and $Q^{abf}$ scores confirm enhanced perceptual quality and edge alignment, demonstrating the method’s robust capability to produce semantically consistent and visually faithful fusion results.
The RoadScene dataset includes complex traffic scenarios with varying illumination and diverse target distributions, imposing high demands on fusion for target detection and scene understanding. Fig. 8 shows visual comparisons. Methods such as U2Fusion and IGNet fail to retain both thermal cues and visual details, leading to blurred textures or weakened targets. Text-DiFuse and SAGE show limited target enhancement and detail clarity. MaeFuse preserves some textures but often produces dark results, reducing visual quality. In contrast, our method achieves superior brightness balance, structural clarity, and background suppression, demonstrating strong robustness and practical value under complex conditions.
4.3 Efficiency Comparison
To comprehensively compare with other state-of-the-art fusion methods, we evaluated the runtime efficiency of each approach in terms of computational cost and parameter scale. The detailed results are presented in Fig. 9. Computational cost is measured by the number of floating-point operations (FLOPs) required to process a single image, using the MSRS dataset for testing. TarDAL achieves higher efficiency due to its lightweight architecture with fewer parameters and lower computational cost; however, its fusion performance in terms of visual quality is inferior to our method. Diffusion-based methods introduce diffusion models, which significantly increase both parameter count and computational burden. SAGE incorporates SAM-generated segmentation masks as prompts to enhance semantic awareness in fusion but at the expense of increased computational complexity. Text-DiFuse employs a text-modulated diffusion model, improving flexibility and representational capacity in fusion, yet this also leads to higher computational overhead. In contrast, our method achieves a favorable balance by maintaining high-quality fusion results with moderate parameter size and computational cost, thus delivering better runtime efficiency. Overall, the analysis demonstrates that our approach strikes an effective trade-off between fusion performance and computational resource consumption, highlighting strong competitiveness and practical applicability.
4.4 Ablation Study
To thoroughly investigate the contribution of each component in the network architecture, a series of ablation studies were carefully designed and conducted, with systematic evaluations performed on the MSRS dataset.
4.4.1 Impact of SAM
The Segment Anything Module (SAM) offers high-quality semantic masks in stage II, guiding the diffusion model for structural refinement and semantic enhancement, thereby improving image integrity and consistency. To assess SAM’s effectiveness, we conducted ablation studies by removing
Table 3: Quantitative results of all fusion variants in the four ablation studies. Bold numbers indicate the best performance.
MSRS EN SD SF MI SCD VIF $Q^{abf}$
I. Impact of SAM
w/o SAM 6.43 43.27 11.98 2.54 1.48 0.87 0.62
w/o IR SAM 6.68 44.73 12.86 2.79 1.69 0.96 0.70
w/o VIS SAM 6.72 44.82 12.79 2.84 1.67 0.98 0.68
II. Impact of Two-Stage Training
w/o Stage I 6.71 45.03 13.03 2.87 1.69 0.98 0.71
w/o Stage II 6.52 44.18 12.87 2.64 1.64 0.94 0.66
III. Impact of the Diffusion Process
w/o Dif 6.69 44.89 12.96 2.83 1.62 0.99 0.63
IV. Impact of HFAH
w/o HFAF 6.74 45.05 13.14 2.91 1.67 1.03 0.68
Ours method 6.81 45.28 13.27 2.99 1.73 1.06 0.74
or modifying its components, designing three variants to analyze the impact of semantic guidance on fusion performance.
(1). Variant (a): Removing the SAM module (w/o SAM)
In this setting, we completely remove the SAM module and replace the semantic masks with randomly cropped patches from the original images, thereby fully eliminating the guidance of semantic priors in the diffusion process. This allows us to evaluate the model’s performance in the absence of high-level semantic information.
(2). Variant (b): Removing the semantic mask of the infrared image (w/o IR SAM)
This variant retains only the semantic mask generated from the visible image while removing the semantic information of the infrared modality, aiming to analyze the specific role of thermal radiation structural features from the infrared image in guiding the optimization process.
(3). Variant (c): Removing the semantic mask of the visible image (w/o VIS SAM)
Contrary to the previous setting, this
Figure 10: Quantitative ablation results on the MSRS dataset regarding the number of repeated MSFEM and TB modules.
variant retains the infrared semantic mask while removing the visible light mask, thereby evaluating the importance of the rich texture and edge information from the visible images in detail preservation and target recognition.
As shown in Table III, variant (a) without SAM performs worst in all metrics, highlighting the importance of semantic priors for structural and semantic fidelity. Variants (b) and (c) show milder drops but remain inferior to the full model, confirming the complementary benefits of dual-modality semantic masks. These results validate SAM’s critical role in enhancing semantic representation and downstream task adaptability.
4.4.2 Analysis of the Impact of the Number of MSFEM and TB Modules
To explore the optimal stacking depth for feature extraction, we performed ablation experiments by varying the number of MSFEM and TB modules in Stage I from 2 to 4 to investigate the influence of stacking depth on fusion performance. As shown in Fig. 10, increasing the number of modules generally improves fusion quality by enhancing detail restoration and structural consistency. However, performance saturates or slightly declines beyond three layers due to potential redundancy and overfitting. Therefore, we adopt three MSFEM and three TB layers as the final configuration, striking a balance between accuracy and efficiency, and providing a strong feature foundation for stage II diffusion optimization.
4.4.3 Impact of Two-Stage Training
The proposed fusion framework adopts a two-stage training strategy: Stage I uses MSFEM and TB for modal feature extraction and preliminary fusion, while Stage II incorporates SAM and a conditional diffusion model for structural refinement under semantic guidance. To evaluate this design, we conducted ablation experiments comparing the full method with variants using only Stage I or Stage II. As shown in Table III, the Stage-I-only model (w/o Stage II) preserves basic structure but lacks detail and semantic consistency, while the Stage-II-only variant (w/o Stage I) yields blurred and distorted results due to missing structural priors. In contrast, the full two-stage strategy synergistically combines low-level structure with high-level semantics, enhancing alignment, detail preservation, and robustness in downstream tasks.
4.4.4 Impact of the Diffusion Process
As the core of the second stage, the diffusion model refines structure and enhances semantics via progressive denoising. To verify its effect, we conducted an ablation by removing the diffusion process (w/o Dif) while keeping the network unchanged. Table III shows that on MSRS, all metrics decline without diffusion, demonstrating its critical role in improving clarity, structural integrity, and fusion quality under semantic guidance. This confirms diffusion modeling’s unique advantage and necessity for high-fidelity, semantically consistent multimodal fusion.
4.4.5 Impact of HFAH
To enhance semantic guidance and structural restoration during denoising, we introduce the Hierarchical Feature Aggregation Head (HFAH) in the decoder of the diffusion model. HFAH aggregates multi-scale features via spatial attention to model edges and semantic regions more effectively. In ablation (w/o HFAH), removing hierarchical fusion and attention leads to noticeable drops in boundary fidelity and texture detail, as shown in Table III—especially in complex scenes. These results confirm that HFAH improves fine-grained structural recovery and semantic consistency.
4.5 Extension to high-level vision task
To validate adaptability in high-level vision tasks, we applied the method to semantic segmentation and object detection, using their performance to comprehensively evaluate fusion quality and demonstrate task-driven advantages.
4.5.1 Object Detection
Object detection experiments on the MSRS dataset using YOLOv5 [39] show that fused images outperform single-modality inputs. Our method achieves the best detection accuracy among mainstream fusion approaches, demonstrating its superior effectiveness and practical value in downstream tasks.
Fig. 11 shows that fused images significantly improve detection accuracy over single-modality inputs. Our method outperforms mainstream
Figure 11: Quantitative evaluation of fusion methods on object detection. Red and blue dots indicate the best and second-best results, respectively.
Figure 12: Visual results of the object detection task on the MSRS dataset.
fusion approaches in both “Person” and “Car” categories, demonstrating superior object perception and generalization.
Table 4: Quantitative comparison of different fusion methods across categories. Best and second-best scores are marked in red and blue, respectively.
Category Background Car Person Curve Color Cone Bike mIoU
U2Fusion 0.976 0.857 0.682 0.703 0.725 0.716 0.776
TarDAL 0.969 0.867 0.681 0.755 0.783 0.769 0.804
PIAFusion 0.972 0.871 0.680 0.721 0.766 0.739 0.791
IGNet 0.971 0.869 0.679 0.769 0.779 0.754 0.804
CDDFuse 0.975 0.872 0.685 0.789 0.794 0.785 0.817
Dif-fusion 0.976 0.871 0.684 0.774 0.772 0.786 0.810
PFCFuse 0.975 0.873 0.683 0.737 0.748 0.729 0.791
SHIP 0.976 0.871 0.681 0.745 0.760 0.742 0.796
EMMA 0.977 0.862 0.656 0.735 0.769 0.638 0.773
Text-Difuse 0.972 0.873 0.673 0.743 0.764 0.748 0.796
DAFusion 0.977 0.868 0.679 0.756 0.778 0.750 0.802
MaeFuse 0.975 0.871 0.685 0.774 0.796 0.787 0.814
SAGE 0.981 0.879 0.675 0.747 0.783 0.774 0.806
Ours 0.979 0.879 0.686 0.780 0.798 0.789 0.819
Figure 13: Visualization results of semantic segmentation. We compare our method with seven classical approaches for qualitative analysis. The regions with significant differences are highlighted in red boxes.
As shown in Fig. 12, our method outperforms others in accurately localizing multiple targets by effectively preserving infrared thermal information and integrating visible details, thereby enhancing detection accuracy, robustness, and adaptability in complex scenes.
4.5.2 Semantic Segmentation
We use DeeplabV3+ [40] for semantic segmentation evaluation, measuring performance with pixel-level IoU. DeeplabV3+ is a well-established model with strong feature modeling, making it suitable to assess how well the fused images support downstream segmentation tasks.
Table V shows segmentation results for six categories (Car, Person, Bike, Color Cone, Background, Curve). Our method achieves the highest IoU in most classes—especially Background, Car, and Person—and significantly outperforms others in mean IoU, demonstrating strong generalization and semantic consistency across categories.
Furthermore, Fig. 13 shows semantic segmentation visualizations on complex scenes. U2Fusion and CDDFuse exhibit blurred boundaries and texture loss, struggling with multiple targets and fine details. MaeFuse and SAGE improve coherence but still missegment small objects and complex backgrounds. Our method achieves clearer boundaries and more complete contours, demonstrating superior structural fidelity and semantic consistency.
5 DISCUSSION
This paper proposes SGDFuse, an innovative framework for infrared and visible image fusion, designed to address the prevalent semantic gap in current fusion methods. Our core contribution lies in establishing a novel fusion paradigm: leveraging prior knowledge from a large segmentation model (SAM) as explicit semantic guidance to constrain and optimize the generative process of a conditional diffusion model, thereby achieving high-fidelity synthesis of cross-modal information. Extensive quantitative experiments, along with evaluations on downstream tasks such as object detection and semantic segmentation, consistently demonstrate that the fused images generated by SGDFuse significantly surpass state-of-the-art methods in terms of information integrity, detail fidelity, and semantic consistency. This robustly validates the superiority and effectiveness of our proposed ”semantic-guided generation” framework.
We have carefully considered the trade-off between performance and efficiency in our framework. The introduction of SAM and a diffusion model in SGDFuse admittedly increases computational complexity. However, we argue that this design is a necessary approach to overcome the fundamental limitations of traditional methods. Because traditional methods cannot accurately distinguish between foreground targets and background textures, they often lead to the loss of critical thermal radiation information or detail distortion, which fundamentally compromises the quality of the fused information. In contrast, SGDFuse exchanges a manageable computational overhead for a substantial leap in the semantic fidelity of the fused information.
Despite the state-of-the-art performance achieved by SGDFuse, there remain avenues for further exploration that constitute our future research directions. First, since SAM’s pre-training data is primarily based on visible light images, its generalization capability to the infrared domain has room for improvement. In the future, the precision of the semantic guidance can be further enhanced through domain-adaptive fine-tuning or by incorporating more specialized infrared segmentation models. Second, the iterative nature of the diffusion model limits its application in real-time scenarios. Future work will focus on model lightweighting, for instance, through knowledge distillation or by designing efficient samplers, to develop a version suitable for real-time applications and thereby broaden its operational boundaries.
6 CONCLUSION
Existing infrared and visible image fusion (IVIF) methods generally face a bottleneck: an inability to deeply understand the deep semantic information of images. To address this, we propose a novel framework named SGDFuse. At the core of our method is a synergistic two-stage design: the first stage is responsible for robust multi-modal feature extraction and preliminary fusion, while the second stage leverages semantic priors from SAM as explicit guidance to drive a conditional diffusion model for high-fidelity reconstruction. This design enables the simultaneous preservation of salient targets with semantic precision and the restoration of fine-grained texture details, effectively unifying the dual objectives of semantic consistency and visual quality. Extensive experiments conducted on public benchmark datasets demonstrate that our method significantly outperforms current state-of-the-art approaches in both subjective and objective evaluations. Furthermore, SGDFuse exhibits superior performance and generalization capability on downstream tasks, including object detection and semantic segmentation. Future work will focus on optimizing the framework for real-time performance and extending this semantic-guided generative paradigm to other multi-modal fusion applications.
Acknowledgments
This research was supported by the National Natural Science Foundation of China (62301105, 61772319, 62272281, 62002200, 62202268), Shandong Natural Science Foundation of China (ZR2020QF012 and ZR2021MF068), Yantai science and technology innovation development plan(2022JCYJ031).
Data Availability Statement
Data available on request from the authors.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.